Dhabitah, a 60 kg diver is positioned so that her radius of gyration is 0.5 m as she leaves the board with an angular velocity of 4 rads-1. Compute Dhabitah’s angular velocity when she assumes a tuck position, altering her radius of gyration to 0.25 m. Explain how the diver applied the conservation of angular momentum to perform a somersault. 


1
Expert's answer
2022-02-08T10:14:54-0500

Answer

Given data are

Mass of diver m=60kg


radius of gyration is K"_1" =0.5 m


an angular velocity of"\omega_1="

1 4 rads-1


Now


radius of gyration is K2=0.25 m


an angular velocity of"\omega_2" = ?


applied the conservation of angular momentum to perform

"I_1\omega_1=I_2\omega_2\\\omega_2=\frac{I_1\omega_1}{I_2}
\\\omega_2=\frac{mK_1^2\omega_1}{mK_2^2}"

So putting all values

"\omega_2=\frac{(0.5)^2\times4}{(0.25)^2}\\=16rad/sec."